OpenStudy (anonymous):

Help please. What is the x-coordinate of the x-intercept of the line that passes through points (-5, 3) and (1, -3)?

OpenStudy (anonymous):

I know that part

OpenStudy (anonymous):

\[\text { To calculate the slope of a line , use the formula}\\ m=\frac{ y _{2}-y_{1} }{x_{2}-x_{1}}\\ \text { Here } y _{2}=\ -3 \ \ \ and y_{1}=3\\\\ x _{2}= 1\ \ \ \text { and } x_1{-5}\\\\m=\frac{-3-3}{1-(-5)} =-1\\\\\text{ Then just substitute the slope In place of m, and the value of the} \\\\ \text{ y intercept In place of the equation y = mx + b } \\\\(3)=(-1)*(-5)+b \\ b=-2\\\\ y=-x-2\]

OpenStudy (anonymous):

So -2 is the correct answer ?

OpenStudy (anonymous):

Ok I thought that was the answer but wasn't sure I was doing it correct, so thank you.